# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ 4.1 Proof of Lemma 1
Since $\left\lVert\eta\right\rVert_{1}=1$, $\eta$ can be seen as a discrete probability measure. Assume $(G_{i})_{i\geq 1}$ is a sequence of i.i.d. random variables with distribution $\mathbb{P}(G_{i}=n)=\eta_{n}$ for $n\geq 1$. We first consider the rescaled random sum
@@
where $I^{T}$ is a geometric random variable with parameter $1-a_{T}$. While $J^{T}$ is a discrete random variable, our goal is to show that its distribution converges to a continuous one with a specific density. To this end, we study the Laplace transform of $J^{T}$, denoted by $\hat{\rho}^{T}$, which for $z\geq 0$ satisfies:
$\hat{\rho}^{T}(z)=$ $\mathbb{E}[e^{-zJ^{T}}]=\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\mathbb{E}\left[e^{-\frac{z}{T}\sum_{i=1}^{k}G_{i}}\right]$ (22)
@@
By the Abel’s lemma, we obtain
$\sum_{n=0}^{\infty}\eta_{n}e^{-zn}=$ $\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=1}^{n}\eta_{k}=\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\left(1-\sum_{k=n+1}^{\infty}\eta_{k}\right)$
$=$ $1-\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=n+1}^{\infty}\eta_{k}=1-(1-e^{-z})\sum_{n=0}^{\infty}e^{-zn}\sum_{k=n+1}^{\infty}\eta_{k}.$
Since $\eta_{k}\sim\frac{K}{k^{\alpha+1}}$ for sufficiently large $k$, it follows that $\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}n^{-\alpha}$. From Lemma 3, we obtain
@@
$v_{T}=\delta^{-1}T^{\alpha}(1-a_{T}),\quad\text{where}\quad\delta=K\frac{\Gamma(1-\alpha)}{\alpha}.$ (23)
Under Assumption 2, this sequence converges to a limit, which we denote by $v_{\infty}$:
$\lim_{T\to\infty}v_{T}=\lim_{T\to\infty}\delta^{-1}T^{\alpha}(1-a_{T})=\delta^{-1}(\lambda\delta)=\lambda=:v_{\infty}.$
Substituting the expansion of $\hat{\gamma}(z/T)$ into the expression (22) for $\hat{\rho}^{T}(z)$ and taking the limit as $T\to\infty$, we find that for any fixed $z\geq 0$:
$\lim_{T\to\infty}\hat{\rho}^{T}(z)=\frac{\lambda}{\lambda+z^{\alpha}}=\frac{v_{\infty}}{v_{\infty}+z^{\alpha}}.$
The right-hand side is the Laplace transform of the probability density function $x\mapsto\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$. By Lévy’s continuity theorem for Laplace transforms, the convergence of the transforms implies the weak convergence of the corresponding probability measures. This completes the proof of the lemma.
4.2 Proof of Lemma 2
The proof consists of two main steps. First, we rigorously show that for a function $f(t):=t^{\alpha}l(t)$, which is regularly varying with index $\alpha>-1$, its sum is asymptotically equivalent to its integral. Second, we apply Karamata’s Integral Theorem.
First, let’s prove that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)dt$. Our goal is to prove that $\lim_{x\to\infty}\left(\sum_{t=X}^{x}f(t)\right)/\left(\int_{X}^{x}f(t)dt\right)=1$. We use the Stolz–Cesàro theorem. Let $a_{x}=\sum_{t=X}^{x}f(t)$ and $b_{x}=\int_{X}^{x}f(t)dt$. Since $\alpha>-1$, the function $f(t)$ is eventually positive, and thus $b_{x}$ is strictly increasing to infinity, satisfying the conditions for the theorem. We therefore analyze the limit of the ratio of differences:
According to the Uniform Convergence Theorem for regularly varying functions (Bingham et al., 1989, Theorem 1.5.2), the convergence of $f(\lambda x)/f(x)\to\lambda^{\alpha}$ is uniform for $\lambda$ in any compact interval $[a,b]$ with $0<a\leq b<\infty$. Let us choose such an interval that contains $1$, for example $[1/2,2]$. For $x$ large enough, our interval [11/x,1][1-1/x,1] is contained within $[1/2,2]$.
$(1-\varepsilon)f(x)\leq f(t)\leq(1+\varepsilon)f(x).$
This demonstrates that $\int_{x-1}^{x}f(t)dt\sim f(x)$ as $x\to\infty$. Therefore, the limit of the ratio of differences is 1. By the Stolz-Cesàro theorem, the limit of the original ratio is also 1, which proves that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)dt$.
Next, we apply Karamata’s Integral Theorem. Furthermore, Karamata’s theorem for integrals of regularly varying functions (Bingham et al., 1989, Proposition 1.5.8) states that
$\int_{X}^{x}t^{\alpha}l(t)dt\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x)\quad\text{as }x\to\infty.$
Combining the above asymptotic equivalences, we obtain the desired result:
$\sum_{t=X}^{x}t^{\alpha}l(t)\sim\int_{X}^{x}t^{\alpha}l(t)dt\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x).$
This completes the proof.
4.3 Proof of Lemma 3
@@
Now, to find the asymptotic behavior of $\hat{V}(s)$, we first need the asymptotics of $V_{n}$. We are given that $\beta_{k}\sim\frac{ck^{\rho}l(k)}{\Gamma(1+\rho)}$. This is precisely the form required by Lemma 2. Applying the lemma, we find the asymptotic behavior of the partial sums:
@@
With the behavior of $V_{n}$ established, we can apply Karamata’s Tauberian Theorem (e.g., Corollary 1.7.3 in Bingham et al. (1989)) to connect the asymptotics of the partial sums to their transform. The theorem yields:
$\hat{V}(s)\sim cs^{-(\rho+1)}l(1/s)\quad\text{as }s\downarrow 0.$
@@
4.4 Proof of Lemma 4
We use the criterion from Lemma 5 to verify the $C$-tightness for the sequence of càdlàg processes $(Y_{t}^{T})_{t\in[0,1]}$.
Before verifying the conditions, we establish the asymptotic orders of the key moments for large $T$ and for any time index $s$ such that $s/T\to t\in(0,1]$. We use the decomposition $\lambda_{s}^{T}=q_{s}^{T}+\kappa_{s}^{T}$, where $q_{s}^{T}$ is the deterministic mean and $\kappa_{s}^{T}$ is the zero-mean component.
  • We first calculate the first Moment of $\lambda_{s}^{T}$. We have
    $\mathbb{E}[\lambda_{s}^{T}]=q_{s}^{T}=\mu^{T}+\mu^{T}\sum_{u=1}^{s-1}A_{u}^{T}\leq\frac{\mu^{T}}{1-a_{T}}\sim\frac{\nu^{*}}{(\lambda\delta^{2})}T^{2\alpha-1}.$
    Hence, there exists a constant $\bar{C}_{1}>0$ such that $\mathbb{E}[\lambda_{s}^{T}]\leq\bar{C}_{1}T^{2\alpha-1}$ for $T\geq 1$.
  • Next, we calculate the second Moment of $\lambda_{s}^{T}$. The second moment is $\mathbb{E}[(\lambda_{s}^{T})^{2}]=(q_{s}^{T})^{2}+\mathbb{E}[(\kappa_{s}^{T})^{2}]$. We now rigorously derive the order of $\mathbb{E}[(\kappa_{s}^{T})^{2}]$. The exact expression is:
    $\mathbb{E}[(\kappa_{s}^{T})^{2}]=\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\mathbb{E}[\lambda_{s-k}^{T}].$
@@ 4.4 Proof of Lemma 4
    First, we calculate the asymptotic order of $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$. From the proof of Lemma 1, we established that the scaled sequence $TA_{k}^{T}/\left\lVert A^{T}\right\rVert_{1}$ converges to a limiting density. For large $k$ and $T$ such that $k/T\to x>0$, this implies:
    $A_{k}^{T}\sim\frac{\left\lVert A^{T}\right\rVert_{1}}{T}\left(\lambda(k/T)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})\right).$
    From Assumption 2, we have the following asymptotics:
    $\left\lVert A^{T}\right\rVert_{1}=\frac{a_{T}}{1-a_{T}}\sim\frac{1}{1-a_{T}}\sim\frac{T^{\alpha}}{\lambda\delta}.$
    Substituting this into the expression for $A_{k}^{T}$:
    $A_{k}^{T}$ $\sim\frac{T^{\alpha}}{\lambda\delta T}\left(\lambda(k/T)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})\right)$
    $=\frac{T^{\alpha-1}}{\delta}\cdot T^{1-\alpha}k^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})$
    $=\frac{1}{\delta}k^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha}).$
    For large $k$, the Mittag-Leffler function term is bounded. This shows that $A_{k}^{T}$ is regularly varying with index $\alpha-1$. Consequently, $(A_{k}^{T})^{2}$ is regularly varying with index $2\alpha-2$.
    To determine the asymptotic order of the sum more rigorously, we first approximate it by its corresponding integral. Using the asymptotic form of $A_{k}^{T}$ derived from the proof of Lemma 1, for $s=tT$, we have:
    $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ $\sim\int_{1}^{s-1}(A_{x}^{T})^{2}dx$
    $\approx\int_{1}^{tT}\left(\frac{1}{\delta}x^{\alpha-1}E_{\alpha,\alpha}(-\lambda(x/T)^{\alpha})\right)^{2}dx$
    $=\frac{1}{\delta^{2}}\int_{1}^{tT}x^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda(x/T)^{\alpha})\right]^{2}dx.$
    To properly evaluate the dependence on $T$, we perform a change of variables by letting $u=x/T$. This implies $x=uT$ and dx=Tdudx=T\,du. The limits of integration transform from $[1,tT]$ to $[1/T,t]$. Substituting these into the integral yields:
    $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ $\approx\frac{1}{\delta^{2}}\int_{1/T}^{t}(uT)^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}(T\,du)$
    $=\frac{T^{2\alpha-1}}{\delta^{2}}\int_{1/T}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du.$
    As $T\to\infty$, the lower limit of the integral $1/T\to 0$. The integral itself converges to a finite, positive constant that depends on $t$ but not on $T$:
    $\lim_{T\to\infty}\int_{1/T}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du=\int_{0}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du=:C_{*}(t).$
    This limit is well-defined and finite because as $u\to 0$, the integrand behaves like $u^{2\alpha-2}$, and since $\alpha>1/2$, the exponent $2\alpha-2>-1$, ensuring the integral converges at the origin. Thus, for any $t\in[0,1]$, we have established the asymptotic order:
    $\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\sim\frac{C_{*}(t)}{\delta^{2}}T^{2\alpha-1},$
    where
    $C_{*}(t)=\int_{0}^{t}u^{2\alpha-2}[E_{\alpha,\alpha}(-\lambda u^{\alpha})]^{2}du.$
    Since the integrand is non-negative, $C_{*}(t)$ is a non-decreasing function of $t$. This implies that for any $t\in[0,1]$, $C_{*}(t)\leq C_{*}(1)$. Therefore, there exists a uniform constant $\bar{C}_{A}:=C_{*}(1)/\delta^{2}$, independent of $t$ and $T$, such that for all sufficiently large $T$:
    $\sup_{t\in[0,1]}\left(T^{-(2\alpha-1)}\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\right)\leq\bar{C}_{A}.$
    This provides the necessary uniform bound for the sum.
@@
    (This uniform constant $\bar{C}_{2}$ is obtained by taking the bound corresponding to $t=1$, since the underlying integral defining the bound is non-decreasing in $t$).
@@
    $\mathbb{E}[(X_{s}^{T})^{2}]\leq\bar{C}_{1}T^{2\alpha-1}+\left((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$ (24)
@@
Now let us verify the conditions for $C$-tightness.
@@
    Condition 2 (Small Increments in Probability): Given $\theta>2$, for large $T$, the sum $\sum_{s=\lfloor kT^{1-\theta}\rfloor+1}^{\lfloor(k+1)T^{1-\theta}\rfloor}X_{s}^{T}$ contains at most one term. This justifies:
    $\sup_{k,h}\left|\Delta_{h}Y_{k/T^{\theta}}^{T}\right|\leq\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sup_{s\in\{1,\dots,\lfloor T\rfloor\}}X_{s}^{T}.$
@@
    From (24), $\mathbb{E}[(X_{i}^{T})^{2}]\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$. Hence,

### v7

@@ Appendix A Proof of Lemmas
We set $\eta_{0}:=0$ and consequently $A_{0}^{T}:=0$ for all $T$. Since $\|\eta\|_{1}=1$ and $\eta_{n}\geq 0$ for all $n\geq 1$, $\eta$ can be seen as a discrete probability measure. Moreover, for each $T$ we have $\eta_{n}^{T}=a_{T}\eta_{n}\geq 0$, so $(\eta_{n}^{T})_{n\geq 1}$ is also a non-negative sequence of reproduction coefficients. Assume $(G_{i})_{i\geq 1}$ is a sequence of i.i.d. random variables with distribution $\mathbb{P}(G_{i}=n)=\eta_{n}$ for $n\geq 1$. We first consider the rescaled random sum
@@
where $I^{T}$ is a geometric random variable with parameter $1-a_{T}$. Set $S^{T}:=\sum_{i=1}^{I^{T}}G_{i}\in\mathbb{N}$. Since $(G_{i})$ are i.i.d. with pmf $(\eta_{n})$, for $n\geq 1$ we have
$\mathbb{P}(S^{T}=n)=\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\,\mathbb{P}\!\left(\sum_{i=1}^{k}G_{i}=n\right)=\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\,(\eta^{*k})_{n}.$}
Recalling that $\eta^{T}=a_{T}\eta$, this can be rewritten as
$\mathbb{P}(S^{T}=n)=\frac{1-a_{T}}{a_{T}}\sum_{k=1}^{\infty}(\eta^{T*k})_{n}.$}
By the definition of the renewal sequence $(A_{n}^{T})$ (cf. (2.1)), we have $A_{n}^{T}=\sum_{k=1}^{\infty}(\eta^{T*k})_{n}$, and since $\sum_{n\geq 1}(\eta^{T*k})_{n}=\|\eta^{T}\|_{1}^{k}=a_{T}^{k}$, it follows that $\|A^{T}\|_{1}=\sum_{k=1}^{\infty}a_{T}^{k}=a_{T}/(1-a_{T})$. Hence
$\mathbb{P}\!\left(J^{T}=\frac{n}{T}\right)=\mathbb{P}(S^{T}=n)=\frac{A_{n}^{T}}{\|A^{T}\|_{1}},\qquad n\geq 1.$}
Note that $J^{T}$ is discrete, whereas $\rho^{T}$ is a step density. To connect them, let $U$ be uniformly distributed on $(0,1)$, independent of everything else, and define the “jittered” variable $\widetilde{J}^{T}:=J^{T}+U/T$. Then $\widetilde{J}^{T}$ is absolutely continuous on $0,\infty)$ and, by construction, it has density $\rho^{T}$ as defined in (2.8).
To identify the limit law, we study the Laplace transform of $J^{T}$ (and thus of $\widetilde{J}^{T}$). Denote $\hat{\rho}^{T}(z):=\mathbb{E}[e^{-zJ^{T}}]$, $z\geq 0$. It satisfies
$\hat{\rho}^{T}(z)=$ $\mathbb{E}[e^{-zJ^{T}}]=\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\mathbb{E}\left[e^{-\frac{z}{T}\sum_{i=1}^{k}G_{i}}\right]$ (A.1)
@@
Moreover, since $\widetilde{J}^{T}=J^{T}+U/T$ with $U\sim\mathrm{Unif}(0,1)$ independent, the Laplace transform of the density $\rho^{T}$ satisfies
$\int_{0}^{\infty}e^{-zx}\rho^{T}(x)\,dx=\mathbb{E}[e^{-z\widetilde{J}^{T}}]=\mathbb{E}[e^{-zJ^{T}}]\,\mathbb{E}[e^{-zU/T}]=\hat{\rho}^{T}(z)\,\frac{1-e^{-z/T}}{z/T}.$}
Since $(1-e^{-z/T})/(z/T)\to 1$ as $T\to\infty$, the limiting Laplace transform for $\rho^{T}$ is the same as that obtained for $\hat{\rho}^{T}(z)$.
By Abel’s lemma, we obtain
${\sum_{n=1}^{\infty}\eta_{n}e^{-zn}}=$ ${\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=1}^{n}\eta_{k}}$
$=$ ${\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\left(1-\sum_{k=n+1}^{\infty}\eta_{k}\right)}$
$=$ ${1-\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=n+1}^{\infty}\eta_{k}}$
$=$ ${1-(1-e^{-z})\sum_{n=0}^{\infty}e^{-zn}\sum_{k=n+1}^{\infty}\eta_{k}.}$
By Assumption 2.1(i) and Lemma 2.2, we have
$\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}\,n^{-\alpha}\quad\text{as}\quad n\to\infty.$}
From Lemma 2.5, we obtain
@@
$v_{T}=\delta^{-1}T^{\alpha}(1-a_{T}),$ (A.2)
where we recall that $\delta=K\frac{\Gamma(1-\alpha)}{\alpha}$ (see Assumption 2.3). Under Assumption 2.3, we have
$\lim_{T\to\infty}v_{T}=\lambda.$}
Substituting the expansion of $\hat{\gamma}(z/T)$ into (A.1) and taking $T\to\infty$, we find that for any fixed $z\geq 0$:
$\lim_{T\to\infty}\hat{\rho}^{T}(z)=\frac{\lambda}{\lambda+z^{\alpha}}.$}
By (2.7), the right-hand side coincides with the Laplace transform of the limiting density $f^{\alpha,\lambda}$ (defined in (2.5)). By Lévy’s continuity theorem for Laplace transforms, the convergence of the transforms implies the weak convergence of the corresponding probability measures.
Since $P^{T}\stackrel{{\scriptstyle\mathrm{w}}}{{\longrightarrow}}P$ and $F^{\alpha,\lambda}$ is continuous on $0,\infty)$, we now justify the uniform convergence of the distribution functions on $[0,1]$. Recall that $F^{T}(t):=\int_{0}^{t}\rho^{T}(x)\,dx$ is the distribution function of $P^{T}$, hence $F^{T}$ is nondecreasing, right-continuous, and satisfies $\lim_{t\downarrow 0}F^{T}(t)=0$ and $\lim_{t\uparrow\infty}F^{T}(t)=1$. Therefore, as a sequence of distribution functions converging weakly to a distribution function with continuous limit $F^{\alpha,\lambda}$, Pólya’s theorem (see e.g. (Chung, 2001, Chapter 4.3, Exercise 4)) yields
$\sup_{t\in[0,1]}\left|F^{T}(t)-F^{\alpha,\lambda}(t)\right|\longrightarrow 0,\qquad T\to\infty.$
which completes the proof.
A.3 Proof of Lemma 2.5
@@
Now, to find the asymptotic behavior of $\hat{V}(s)$, we first need the asymptotics of $V_{n}$. We are given that $\beta_{k}\sim\frac{ck^{\rho}l(k)}{\Gamma(1+\rho)}$. This is precisely the form required by Lemma 2.2. Applying the lemma, we find the asymptotic behavior of the partial sums:
@@
With the behavior of $V_{n}$ established, we apply Karamata’s Tauberian theorem for power series (e.g., (Bingham et al., 1989, Corollary 1.7.3)). Since
$V_{n}\sim\frac{c}{\Gamma(\rho+2)}\,n^{\rho+1}l(n),$
the Tauberian theorem yields
$\hat{V}(s)\sim\frac{c}{\Gamma(\rho+2)}\,\Gamma(\rho+2)\,s^{-(\rho+1)}l(1/s)=c\,s^{-(\rho+1)}l(1/s)\quad\text{as }s\downarrow 0.$}
@@
A.4 Proof of Lemma 2.6
We use the criterion from Lemma 2.8 to verify the $C$-tightness for the sequence of càdlàg processes $(Y_{t}^{T})_{t\in[0,1]}$.
Before verifying the conditions, we derive moment bounds that hold uniformly in the time index $s\in\{1,\dots,\lfloor T\rfloor\}$. When convenient, we also consider the macroscopic scaling $s=\lfloor tT\rfloor$ with $t\in(0,1]$. We decompose
$\lambda_{s}^{T}=q_{s}^{T}+\kappa_{s}^{T},\qquad q_{s}^{T}:=\mathbb{E}[\lambda_{s}^{T}],\qquad\kappa_{s}^{T}:=\lambda_{s}^{T}-q_{s}^{T},$}
so that $\mathbb{E}[\kappa_{s}^{T}]=0$.
  • We first calculate the first Moment of $\lambda_{s}^{T}$. By Remark 2.12, for all $s\geq 1$,
    $\mathbb{E}[\lambda_{s}^{T}]=q_{s}^{T}=\mu^{T}+\mu^{T}\sum_{u=1}^{s-1}A_{u}^{T}.$
    Since $\sum_{u\geq 1}A_{u}^{T}=\|A^{T}\|_{1}=\frac{a_{T}}{1-a_{T}}$, it follows that
    $q_{s}^{T}\leq\mu^{T}\left(1+\sum_{u\geq 1}A_{u}^{T}\right)=\frac{\mu^{T}}{1-a_{T}}.$
    Consequently, using Assumption 2.3, there exists a constant $\bar{C}_{1}>0$ such that
    $\sup_{1\leq s\leq\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]\leq\bar{C}_{1}\,T^{2\alpha-1},\qquad T\geq 1.$ (A.3)
  • Next, we calculate the second Moment of $\lambda_{s}^{T}$. The second moment is $\mathbb{E}[(\lambda_{s}^{T})^{2}]=(q_{s}^{T})^{2}+\mathbb{E}[(\kappa_{s}^{T})^{2}]$. We now rigorously derive the order of $\mathbb{E}[(\kappa_{s}^{T})^{2}]$. Recall that $\kappa_{s}^{T}=\lambda_{s}^{T}-q_{s}^{T}$ is the centered part of the intensity. Using the recursion for $\lambda_{s}^{T}$ and the fact that $q_{s}^{T}=\mathbb{E}[\lambda_{s}^{T}]$, we can write $\kappa_{s}^{T}$ as a moving-average of centered innovations:
    $\kappa_{s}^{T}=\sum_{k=1}^{s-1}A_{k}^{T}\,\xi_{s-k}^{T},\qquad\xi_{j}^{T}:=X_{j}^{T}-\lambda_{j}^{T}.$
    Conditionally on $\mathcal{F}_{j-1}^{T}$, we have $\mathbb{E}[\xi_{j}^{T}\mid\mathcal{F}_{j-1}^{T}]=0$ and $\mathrm{Var}(\xi_{j}^{T}\mid\mathcal{F}_{j-1}^{T})=\mathrm{Var}(X_{j}^{T}\mid\mathcal{F}_{j-1}^{T})=\lambda_{j}^{T}$ (since $X_{j}^{T}\mid\mathcal{F}_{j-1}^{T}\sim\mathrm{Pois}(\lambda_{j}^{T})$). Moreover, $(\xi_{j}^{T})$ is a martingale difference sequence with respect to $(\mathcal{F}_{j}^{T})$. Hence for $k\neq\ell$ (say $k>\ell$), by the tower property,
    $\mathbb{E}[\xi_{s-k}^{T}\,\xi_{s-\ell}^{T}]=\mathbb{E}\!\left[\xi_{s-\ell}^{T}\,\mathbb{E}[\xi_{s-k}^{T}\mid\mathcal{F}_{s-\ell}^{T}]\right]=0,$
    since $\mathbb{E}[\xi_{s-k}^{T}\mid\mathcal{F}_{s-k-1}^{T}]=0$ and $\mathcal{F}_{s-\ell}^{T}\subseteq\mathcal{F}_{s-k-1}^{T}$ when $k>\ell$. Furthermore, using $\mathbb{E}[\xi_{j}^{T}\mid\mathcal{F}_{j-1}^{T}]=0$, we have
    $\mathbb{E}[(\xi_{j}^{T})^{2}\mid\mathcal{F}_{j-1}^{T}]=\mathrm{Var}(\xi_{j}^{T}\mid\mathcal{F}_{j-1}^{T})=\mathrm{Var}(X_{j}^{T}\mid\mathcal{F}_{j-1}^{T})=\lambda_{j}^{T},$
    and therefore $\mathbb{E}[(\xi_{j}^{T})^{2}]=\mathbb{E}[\lambda_{j}^{T}]$. Consequently,
    $\mathbb{E}[(\kappa_{s}^{T})^{2}]=\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\,\mathbb{E}[(\xi_{s-k}^{T})^{2}]=\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\,\mathbb{E}[\lambda_{s-k}^{T}].$
@@
    First, we calculate an upper bound for $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ that is uniform in $s\leq T$.
    Define the (finite) measure on $(0,\infty)$
    $\widetilde{\mu}_{T}(dx):=\frac{T}{\|A^{T}\|_{1}}\sum_{k\geq 1}A_{k}^{T}\,\delta_{k/T}(dx).$
    Lemma 2.4 identifies the limit of the Laplace transforms $\int_{0}^{\infty}e^{-zx}\widetilde{\mu}_{T}(dx)$, which implies $\widetilde{\mu}_{T}\Rightarrow\widetilde{\mu}$ in the sense of weak convergence of finite measures. In particular, for any bounded continuous test function $\varphi$,
    $\int_{0}^{\infty}\varphi(x)\,\widetilde{\mu}_{T}(dx)\longrightarrow\int_{0}^{\infty}\varphi(x)\,\widetilde{\mu}(dx).$
    We now bound $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ uniformly for $s\leq T$. By Lemma 2.7, we have the deterministic pointwise bound $A_{k}^{T}\leq Ck^{\alpha-1}$ for all $k\geq 1$ and all $T$. Hence, for $s\leq T$,
    $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ $\leq C^{2}\sum_{k=1}^{s-1}k^{2\alpha-2}\leq C^{2}\sum_{k=1}^{T}k^{2\alpha-2}\leq C^{2}\left(1+\int_{1}^{T}x^{2\alpha-2}\,dx\right)$
    $=C^{2}\left(1+\frac{T^{2\alpha-1}-1}{2\alpha-1}\right)\leq\bar{C}_{2}\,T^{2\alpha-1},$
    for some constant $\bar{C}_{2}>0$ independent of $T$.
@@
    (In the previous paragraph we obtained $\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\leq C_{A}^{2}\,T^{2\alpha-1}\int_{0}^{1}\frac{u^{2\alpha-2}}{(1+u^{\alpha})^{2}}\,du$; hence we may take $\bar{C}_{2}:=C_{A}^{2}\int_{0}^{1}\frac{u^{2\alpha-2}}{(1+u^{\alpha})^{2}}\,du$, which is independent of $t$ and $T$.)
@@
    $\mathbb{E}[(X_{s}^{T})^{2}]\leq\bar{C}_{1}T^{2\alpha-1}+\left((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$ (A.4)
@@
Now let us verify the three conditions in Lemma 2.8, which provide a sufficient criterion for $C$-tightness of the sequence $(Y^{T})_{T\geq 1}$.
@@
    Condition 2 (Small increments in probability): Fix $\theta>2$ and consider $t=k/T^{\theta}$ and $h\in[0,T^{-\theta}]$ as in Lemma 2.8. Since $Th\leq T^{1-\theta}<1$ for all $T\geq 2$, we have the deterministic floor-function bound
    $0\leq\lfloor T(t+h)\rfloor-\lfloor Tt\rfloor\leq 1.$}
    Hence the increment interval $(\lfloor Tt\rfloor,\lfloor T(t+h)\rfloor]$ contains either no integer or exactly one integer, and therefore
    ${\Delta_{h}Y_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=\lfloor Tt\rfloor+1}^{\lfloor T(t+h)\rfloor}X_{s}^{T}}\quad{\text{is either }0\text{ or equals }\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\,X_{\lfloor T(t+h)\rfloor}^{T}.}$
    In particular, uniformly over $k\in\{0,1,\dots,\lfloor T^{\theta}\rfloor\}$ and $h\in[0,T^{-\theta}]$,
    $\sup_{k,h}\left|\Delta_{h}Y_{k/T^{\theta}}^{T}\right|\leq\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\,\sup_{1\leq s\leq\lfloor T\rfloor}X_{s}^{T}.$}
@@
    From (A.4), $\mathbb{E}[(X_{i}^{T})^{2}]\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$. Hence,